OpenStudy (anonymous):

list the intercepts of the equation y=3x^3-192

OpenStudy (chihiroasleaf):

the intercepts is the value of x when y = 0 so, to find the intercepts you have to solve 3x^3-192 = 0 no, solve this equation factor this equation, if it's possible

OpenStudy (anonymous):

x =

OpenStudy (chihiroasleaf):

3x^3 -192 = 0 3 ( x^3 - 64) = 0 x^3 - 64 = 0 x^3 = 64 now, can you find x?

OpenStudy (anonymous):

the cube is messing me up.

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

not what

OpenStudy (anonymous):

??

OpenStudy (chihiroasleaf):

that's all, so the intercept is point (4,0)

OpenStudy (anonymous):

coudlnt -4,0 work also?

OpenStudy (chihiroasleaf):

no, (-4)^3 = -64

OpenStudy (anonymous):

thank you. can you help with one more?

OpenStudy (chihiroasleaf):

you're welcome... If I could